### v1

# Name Tagging Under Domain Shift via Metric Learning for Life Sciences

Abstract
Name tagging is a key component of Information Extraction (IE), particularly in scientific domains such as biomedicine and chemistry, where large language models (LLMs), e.g., ChatGPT, fall short. We investigate the applicability of transfer learning for enhancing a name tagging model trained in the biomedical domain (the source domain) to be used in the chemical domain (the target domain). A common practice for training such a model in a few-shot learning setting is to pretrain the model on the labeled source data, and then, to finetune it on a hand-full of labeled target examples. In our experiments we observed that such a model is prone to mis-labeling the source entities, which can often appear in the text, as the target entities. To alleviate this problem, we propose a model to transfer the knowledge from the source domain to the target domain, however, at the same time, to project the source entities and target entities into separate regions of the feature space. This diminishes the risk of mis-labeling the source entities as the target entities. Our model consists of two stages: 1) entity grouping in the source domain, which incorporates knowledge from annotated events to establish relations between entities, and 2) entity discrimination in the target domain, which relies on pseudo labeling and contrastive learning to enhance discrimination between the entities in the two domains. We carry out our extensive experiments across three source and three target datasets, and demonstrate that our method outperforms the baselines, in some scenarios by 5% absolute value.
Name Tagging Under Domain Shift via Metric Learning
for Life Sciences
[…]
1 Introduction
Name tagging is a crucial step in IE tasks. Existing models have achieved remarkable performance in the general domain Lin et al. (2020); Wang et al. (2021a); Zhang et al. (2023); Shen et al. (2023b). However, in the scientific domains, e.g., medical or chemical domains, these models usually struggle due to the extremely large quantity of concepts, the wide presence of multi-token entities, and the ambiguity in detecting entity boundaries.
Figure 1: A test example in the chemical domain. The words marked with blue indicators are the chemical entities, and the words marked with red and orange indicators are biomedical entities. The entities in red are mis-labeled by a few-shot model as chemical entities.
Large language models (LLMs) show an impressive performance on various NLP tasks such as question answering or text summarization OpenAI (2022). Models such as ChatGPT OpenAI (2022) can achieve outstanding results given just a few training examples Wang et al. (2022). However, Kandpal et al. (2023) recently report that the performance of these models is proportional to the number of relevant documents present in their pretraining corpus. Thus, one can expect that their performance fluctuates across domains. This is particularly expected to occur across certain scientific subjects, where the data may be scarce. Given the already existing challenges of the name tagging task in the scientific domain—mentioned earlier—this factor can further exacerbate the problem. For instance, in our early few-shot learning experiments, we observed that the results of ChatGPT in the chemical domain name tagging task are significantly worse than those in the general domain. [^1] We report this complementary experiment in Appendix A.
In the present work, we employ transfer learning (Pan and Yang, 2010) to alleviate this problem. Transfer learning methods exploit the label data from one domain (called the source domain) to minimize the prediction error in another domain (namely the target domain). We particularly focus on a realistic setting, where given a large set of labeled data from the source domain and a small set of labeled data from the target domain, the goal is to develop a model for the target domain. [^2] In the literature, this setup is often called the semi-supervised transfer learning setting (Saito et al., 2019). There are more resources in the biomedical domain than in the chemical domain due to funding priorities and BioNLP workshops. Therefore, as a case study, we take the biomedical domain as the source and the chemical domain as the target.
Figure 1 shows the challenges a name tagging model can face in the chemical domain. The model is trained in a few-shot learning setting. Thus, it is trained on labeled biomedical data, and then, further finetuned on a few labeled examples from the chemical domain. The task is to recognize only the chemical entities—ignoring other types of entities. Blue indicators represent chemical entities, while red and orange indicators denote biomedical entities. The model categorizes the entities marked with blue and red as chemical entities. The first observation is that, to an inexpert human, it is difficult to perform the task because the entities are highly domain-specific. The second observation is that the model wrongfully labels some entities from the source domain as the target entities. Therefore, engineers developing such a model face a dilemma: while not using the source data dramatically deteriorates performance, [^3] We empirically support this in the analysis section. simply pretraining with the source data increases the false positive rate. The third, and perhaps the most important, observation is that the examples from the source and target domains can co-occur in the same document. This problem setting contradicts the regular transfer learning setting, where the examples from the source and the target domains are fully disjoint (Ben-David et al., 2010). The latter property poses difficulties in the applicability of the traditional transfer learning methods in this setting.
Our core idea is to train a name tagging model such that it is able to project the representations of the source and target entities into separate regions of the feature space. Such a model can potentially transfer knowledge from the source domain to the target domain by constructing a shared feature space between the two domains. Furthermore, it reduces the similarity between the representations of the entities in the two domains, and consequently, it can potentially minimize the number of source entities that are mislabeled as target entities. To achieve this, our model consists of a pretraining stage on the source data, and a finetuning stage on the target data.
In the pretraining stage, we propose two methods to enrich the feature space with auxiliary data. The auxiliary data is extracted from the event mentions that the entities participate in. Additionally, during this stage, we propose to employ the multi-similarity loss term (Wang et al., 2019), which enables us to partition (or group) the source entities. Our empirical analysis shows that constructing such a feature space during the pretraining stage facilitates our projection step during the finetuning stage. In the finetuning stage, we detect the potentially false positive entities by pseudo-labeling them. Then, we aim to construct a feature space that projects the pseudo-labeled entities and the target entities into separate regions. We achieve this by employing the multi-similarity loss again. We evaluate our method across twelve use cases and show it outperforms the baselines in most experiments, with improvements of up to 5% in absolute value. We also empirically analyze our method and show that each proposed technique is individually effective. Our contributions are threefold:
  • We propose a new pretraining algorithm for the name tagging task in the transfer learning setting. Our algorithm consists of two steps. First, extracting auxiliary information about the source entities through the event mentions they participate in. Second, proposing an entity grouping technique using the multi-similarity loss. Our methods have proven effective for the name tagging task in the target domain. Our study is carried out in the scientific domain, particularly from the biomedical data as the source domain to the chemical data as the target domain. As we stated before, this is a crucial and challenging real world problem. All of our claims, here and later, are only about this task.
  • We propose a finetuning algorithm for the name tagging task in the transfer learning setting. Our algorithm aims to project the target entities and the entities that may be potentially mislabeled into separate regions of the feature space. It consists of two steps. First, detecting the potentially out-of-domain entities by pseudo-labeling them. Second, obtaining dissimilar representations for the two sets of entities using the multi-similarity loss.
  • We conduct extensive experiments across twelve use cases, demonstrating that our method significantly outperforms the baseline models and shedding light on various aspects of our model.
Figure 2: Overview of proposed entity grouping and entity discrimination frameworks. Entity grouping on the source domain is shown in the upper part. Based on event annotations, a set of event embeddings is constructed under two paradigms. Afterward, pairwise auxiliary similarity scores are calculated according to argument embeddings. Extended multi-similarity loss concerning four types of similarities, combined with cross-entropy loss, are jointly learned during pretraining. Entity discrimination on the target domain is shown in the lower part. Pseudo labels are formed by the name tagging model pretrained in the source domain, and in contrast to annotated labels, a multi-similarity loss is injected into finetuning.
2 Background
2.1 Named Entity Recognition
We view the name tagging task as a sequence labeling problem. We denote the training data by $D{=}\left\{(\mathbf{X}_{i},\mathbf{Y}_{i})\right\}_{i=1}^{n}$, where $n$ is the number of input passages (or texts). To train classifier parameter $\theta$, we minimize the loss as follows:
$\mathcal{L}_{NER}=\mathbb{E}_{(\mathbf{X}_{i},\mathbf{Y}_{i}){\sim}D}\left[CE\left(f(\mathbf{Y}_{i}|\mathbf{X}_{i};\theta),\mathbf{Y}_{i}\right)\right],$ (1)
where $CE$ is the cross-entropy loss.
2.2 Transfer Learning
We are given examples from the source domain $\mathbb{S}$ and the target domain $\mathbb{T}$, where the training set size in the target domain is much smaller than the source domain, i.e., $|\mathbf{X}^{\mathbb{T}}|<<|\mathbf{X}^{\mathbb{S}}|$. Given this data, we aim to develop a model for the target domain and minimize its prediction error.
We focus on the name tagging task, and take the biomedical domain as the source and the chemical domain as the target. Note that the data distributions in the two domains are different. Therefore, a model solely trained on the source data is usually not as competitive as one trained on the target data. The baseline solution in this setting, direct transfer, is to pretrain the model on the labeled source data and finetune it on the labeled target data.
2.3 Multi-Similarity Loss
We enhance our name tagging model in the source domain by capturing the similarities between entity pairs. To achieve this, we employ an objective term called the multi-similarity (MS) contrastive loss proposed by Wang et al. (2019) for metric learning. MS can incorporate self-similarity, relative positive similarity, and relative negative similarity. Self-similarity depends on the properties of the data point itself, such as hardness, while negative and positive similarities are measured with respect to an anchor data point.
In the following sections, we use the final encoder hidden states of input tokens as entity representations. If an entity consists of multiple tokens, we take the average of their representations. Additionally, we use the cosine metric to measure the relative similarity between entity pairs.
The multi-similarity loss is calculated in two stages. First, given the i-th entity denoted by $x_{i}$ and its label denoted by $y_{i}$, we aim to extract the most difficult positive and negative entities. This is achieved by thresholding over the relative similarity scores as follows:
$\mathcal{P}_{i}=\{x_{j}|S_{ij}^{+}$ $<\max_{y_{k}\neq y_{i}}S_{ik}+\epsilon\},$ (2)
$\mathcal{N}_{i}=\{x_{j}|S_{ij}^{-}$ $>\min_{y_{k}=y_{i}}S_{ik}-\epsilon\},$
where $\mathcal{P}_{i}$ is the set of positive, $\mathcal{N}_{i}$ is the set of negative, and $\epsilon$ is a margin penalty.
Second, a soft weight score is calculated for the extracted pairs to reflect their importance in the final loss term:
$w_{ij}^{+}$ $=\frac{e^{-\alpha(S_{ij}-\gamma)}}{1+\sum_{k\in\mathcal{P}_{i}}e^{-\alpha(S_{ik}-\gamma)}},$ (3)
$w_{ij}^{-}$ $=\frac{e^{\beta(S_{ij}-\gamma)}}{1+\sum_{k\in\mathcal{N}_{i}}e^{\beta(S_{ik}-\gamma)}},$
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters. We observe that in each set, the weights are the ratio of the self-similarity scores to the sum of all the relative scores in the set.
The final multi-similarity loss is:
$\mathcal{L}_{MS}=$ $\frac{1}{n_{e}}\sum_{i}^{n_{e}}\left\{\frac{1}{\alpha}\log[1+\sum_{k\in\mathcal{P}_{i}}e^{-\alpha(S_{ik}-\gamma)}]\right.$ (4)
$+\left.\frac{1}{\beta}\log[1+\sum_{k\in\mathcal{N}_{i}}e^{\beta(S_{ik}-\gamma)}]\right\},$
where $n_{e}$ is the total number of the entities. Note that the values of $S_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ already contain the weights computed in Equations 3.
3 Proposed Method
Figure 2 illustrates our framework. It consists of two stages, pretraining on the source data, and then, finetuning on the target data. In the source domain, we employ external knowledge to construct an event feature space based on their arguments, and to calculate the auxiliary similarity scores between entities. Then, the similarity scores are used in the multi-similarity loss to shape the entity feature space. In the target domain, we aim to enhance the model ability to distinguish the target entities from the entities that are likely to be mislabeled, such as the entities that potentially belong to the source domain. To achieve this, we propose an algorithm to extract pseudo-labels and employ the multi-similarity loss for the second time.
3.1 Source Domain Pretraining
Using external knowledge, we enrich entity representations for source domain pretraining. Below, we propose two separate approaches to obtain the auxiliary embedding vectors, both exploiting event mentions that the entities appear in. Given the auxiliary vectors, next, we describe our method for calculating the auxiliary similarity scores. Finally, we provide an overview of the pretraining loss function, which incorporates the entity similarity scores and the auxiliary similarity scores.
Concatenation based Event Embedding.
Our first approach relies on an off-the-shelf token encoder pretrained on biomedical data, called SapBERT Liu et al. (2021a). Given an event mention, we encode its arguments using SapBERT, and then concatenate the resulting vectors to obtain the event representation. Note that in some cases an argument may be a nested event, or an event may have varying number of arguments. In those cases, we use vector averaging to compress the representations or padding to fill in the extra argument slots [^4] The details can be found in Appendix B.1..
Sentence-Encoder based Event Embedding.
In our second approach, we use templates generated by a LLM. We begin by extracting all event types from the source domain. We then submit each type and its arguments to the LLM, using prompts to construct a template. A few examples of such templates are reported in Table 1, and a larger set of templates along our prompt instruction can be found in Appendix B.2. Then, for the event mentions in the source data, we complete their corresponding templates by replacing the placeholders with the actual arguments. The resulting passages are sent through an off-the-shelf sentence encoder to obtain the final representation vectors. In our experiments, we use ChatGPT OpenAI (2022) as the LLM, and use the S-PubMed-BERT Deka et al. (2022) as the sentence encoder.
In the result section, we individually evaluate each one of the methods for extracting the auxiliary embedding vectors.
Type
Template
Phosphorylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the addition of a phosphate group at a particular site <Site>, facilitated by another molecule <Cause>.
Acetylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of an acetyl group at a particular site <Site>, catalyzed by another molecule <Cause>.
Pathway
Indicated by the given trigger <Trigger>, involving one or more molecules <Participant> that collaborate to accomplish a specific biological function or response.
Table 1: Examples of templates for sentence-encoder based embeddings. Angle brackets <·> in templates are placeholders to be replaced by actual entities as corresponding arguments.
Auxiliary Similarity.
Given $E(x_{i})$ and $E(x_{j})$ as the sets of auxiliary vectors for the entities $x_{i}$ and $x_{j}$ respectively, we define the similarity between the two entities as the maximum inter-similarity between all the vector pairs across the two sets. More specifically, we formulate it as follows:
$\kappa_{ij}=\max_{\mu\in E(x_{i}),\nu\in E(x_{j})}\frac{\mu^{T}\cdot\nu}{|\mu|\cdot|\nu|}.$ (5)
The value of $\kappa_{ij}$ captures the relatedness between the contexts that the two entities appear in. If $E(x_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}})$ is empty, then we set $\kappa_{i\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}{=}0$.
Contrastive Grouping.
We adapt the multi-similarity loss Wang et al. (2019) to consider the primary similarity scores between entities, which is the cosine similarity between the encoder outputs for the entities, as well as the auxiliary similarity scores described earlier in this section. Our core idea is to assign a higher weight to the more informative pairs. In the case of positive pairs, this translates into assigning a higher weight to the instances that have a smaller primary similarity and a higher auxiliary similarity. In the case of negative pairs, it is the reverse, i.e., assigning a higher weight to the pairs with higher primary similarity and a lower auxiliary similarity.
The intuition behind these design choices is as follows. In the case of positive pairs, a low primary similarity and a high auxiliary similarity potentially mean that the encoder is unable to properly project the entities, but there is a strong external signal that the pair must have similar representations. In the case of negative pairs, a high primary similarity and a low auxiliary similarity potentially mean that the model needs to revise the parameters to take into account the external signal.
To implement these ideas, we exploit the soft weights discussed in Equations 3 to derive the weights for the positive pairs as follows:
$\hat{w}_{ij}^{+}$ $=\frac{1}{e^{-I_{ij}^{+}}+\sum_{k\in\mathcal{P}_{i}}e^{-J_{ik}^{+}+J_{ij}^{+}}},$ (6)
$I_{ij}^{+}$ $=\alpha(\gamma-S_{ij})+\rho\kappa_{ij},$
$J_{ij}^{+}$ $=\alpha S_{ij}-\rho\kappa_{ij}.$
The value of $\hat{w}_{ij}^{+}$ is the second formulation of the soft weights introduced by Wang et al. (2019). We see that instead of only relying on the values of $S_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ to define $I_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}^{+}$, we incorporate the auxiliary similarity scores $\kappa_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ via the hyperparameter $\rho$.
Similarly, we re-define the soft weights for the negative pairs as follows:
$\hat{w}_{ij}^{-}$ $=\frac{1}{e^{I_{ij}^{-}}+\sum_{k\in\mathcal{N}_{i}}e^{J_{ik}^{-}-J_{ij}^{-}}},$ (7)
$I_{ij}^{-}$ $=\beta(\gamma-S_{ij})+\tau\kappa_{ij},$
$J_{ij}^{-}$ $=\beta S_{ij}-\tau\kappa_{ij},$
where $\tau$ is a hyperparameter to balance the contributions of $S_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$ and $\kappa_{\mathord{\color[rgb]{0.67,0.67,0.67}\bullet}}$.
Given the re-defined soft weights, the refined multi-similarity objective (RMS) can be re-written as follows:
$\mathcal{L}_{RMS}=$ $\frac{1}{n_{e}}\sum_{i}^{n_{e}}\left\{\frac{1}{\alpha}\log[1+\sum_{k\in\mathcal{P}_{i}}e^{-\alpha(S_{ik}-\gamma)+\rho\kappa_{ik}}]\right.$ (8)
$+\left.\frac{1}{\beta}\log[1+\sum_{k\in\mathcal{N}_{i}}e^{\beta(S_{ik}-\gamma)-\tau\kappa_{ik}}]\right\},$
where, as mentioned earlier, $\rho$ and $\tau$ are to maintain a balance between the primary and the auxiliary representations. Note that we use the information extracted from events to construct the auxiliary representations. However, additional source of information can be considered if it is present.
The pretraining objective function consists of the supervised name tagging term and the unsupervised RMS term, as follows:
$\mathcal{L}=\mathcal{L}_{NER}+\lambda_{\mathbb{S}}\mathcal{L}_{RMS},$ (9)
where $\lambda_{\mathbb{S}}$ is a penalty term.
3.2 Target Domain Finetuning
The source and target domains in our problem setting share the same context. In the same documents (or even sentences) that the entities from one domain appear, the entities from the other domain may be used too. This makes the recognition task particularly challenging in the target domain for two reasons: the training data in this domain is scarce, and the presence of entities from the source domain can potentially lead to a high false positive rate. Our core idea is that, while finetuning the model on the target data, we train the encoder such that it projects the target entities and the entities that potentially belong to the source domain into separate regions of the feature space. To implement this idea, we employ pseudo labeling along the multi-similarity loss–introduced in §2.3.
Pseudo Labeling.
Given a passage with annotated target entities in the target training data, we use the model introduced in §3.1 to automatically detect the entities that may belong to the source domain. These entities act as pseudo labels in our algorithm. Note that while there may be passages that do not contain such entities, in general, due to the nature of the two domains that we are studying (i.e., biomedical and chemical domains), this is an expected observation. In the results section, we will also empirically support our argument.
Contrastive Discrimination.
In the next step, we enable the model to discriminate between the target and pseudo-labeled source entities. For this purpose, we use the multi-similarity loss. We use multi-similarity loss in Eq. 4 to calculate contrastive objective by defining the labels as follows:
$y_{i}=\left\{\begin{array}[]{lr}0,&x_{i}\in\mathcal{Q}\\ 1,&x_{i}\notin\mathcal{Q}\end{array}\right.$ (10)
where $x_{i}$ represents the entity and $\mathcal{Q}$ is the set of entities with pseudo labels.
The final target domain fine-tuning objective is:
$\mathcal{L}=\mathcal{L}_{NER}+\lambda_{\mathbb{T}}\mathcal{L}_{MS},$ (11)
where $\lambda_{\mathbb{T}}$ is a penalty terms.
4 Experiments
4.1 Experimental Setup
Dataset
# Train/ Valid/ Test
# Train/ Valid/ Test (All)
PC
260 / 89 / 175
ID
150 / 46 / 117
CG
300 / 100 / 200
CHEMDNER
81 / 72 / 2478
2916 / 2907 / 2478
BC5CDR
86 / 88 / 500
500 / 500 / 500
DrugProt
93 / 88 / 600
597 / 597 / 600
Table 2: Overview of the datasets. The top three are the biomedical source datasets, and the bottom three are the chemical target datasets. The target datasets were down-sampled randomly to be used in the few-shot setting.
Target Tasks
CHEMDNER BC5CDR DrugProt
Metrics
Precision
Recall
F1
Precision
Recall
F1
Precision
Recall
F1
Target Only
42.73
51.69
46.77
72.44
85.86
78.51
63.80
67.42
65.49
Direct Transfer
44.11
48.60
46.18
71.92
86.54
78.55
66.36
70.13
68.17
EG(concat)-Only
43.63
51.06
47.03
72.10
87.12
78.89
68.49
66.54
67.49
EG(sentEnc)-Only
43.09
51.60
46.91
73.96
86.56
79.76
66.02
69.97
67.93
ED-Only
45.71
50.13
47.78
75.01
87.80
80.88
70.15 71.79 70.96
EG(concat)+ED
43.83
52.17
47.64
73.70
85.50
79.15
68.01
69.06
68.52
EG(sentEnc)+ED
45.28
51.68
48.26 76.06
86.58
80.95
67.16
69.04
68.08
Table 3: Evaluation results precision, recall, and F1(%) scores on three target tasks with Biomedical Multi-task as source task. All the reported scores are averaged over 3 different random seeds. We include two baselines, along with our methods EG (Entity Grouping), ED (Entity Discrimination), and their combination.
Datasets. As the source datasets, we use three benchmarks: Pathway Curation (PC), Cancer Genetics (CG), and Infectious Diseases (ID). The first two datasets were released by BioNLP Shared Task 2013 Nédellec et al. (2013), and the third one was released by BioNLP Shared Task 2011 Pyysalo et al. (2011). All three datasets have the same format. We aggregate them to create a fourth dataset called the biomedical multi-task dataset. As the target datasets, we use three benchmarks: CHEMDNER Krallinger et al. (2015), BC5CDR Kim et al. (2015), and DrugPort Miranda et al. (2021).
Each dataset contains extra annotations unrelated to its domain. For instance, the CG dataset has annotations not relevant to biomedicine.
We pre-process all the datasets by removing such annotations.
For few-shot experiments, we down-sample the training and validation sets of target datasets to sizes randomly chosen between 70 and 100. Table 2 summarizes the dataset statistics.
Baselines.
We compare our method with two baselines. Target Only: a model finetuned on the labeled target data.
Direct Transfer: a model pretrained on the labeled source data and then finetuned on the labeled target data.
Implementation Details.
We use BERT (base-uncased) Devlin et al. (2019) as the backbone for all the models. To train the model, we only update the parameters of the adapter layers Houlsby et al. (2019) and freeze the rest. We iteratively select each source and target dataset pair as the training and evaluation benchmarks. All the experiments are repeated three times. Following Nakayama (2018), we report average macro Precision, Recall, and F1 scores [^5] Complementary information about training and tuning can be found in Appendix C..
4.2 Main Results
[…]
Target Tasks
CHEMD BC5CDR DrugProt
Direct Transfer
46.18
78.55
68.17
Pse-Augment
47.14
76.43
68.20
Pse-Classifier
46.93
78.80
69.42
Ours
47.78 80.88 70.96
Table 4: F1(%) scores of our method compared to two alternative methods for using pseudo-labels.
4.3 Empirical Analysis
Pseudo Label Usage. Table 4 reports a comparison between our model and two alternative methods. Pse-Augment, where the detected pseudo-labels are marked and augmented with the target entities and a classifier trained to label unseen target entities. Pse-Classifier, where a classifier is trained to detect pseudo-labels and to filter them out, before being potentially mislabeled. This experiment aims to reveal the efficacy of the multi-similarity (MS) loss for discriminating between the source and the pseudo-labeled entities. The results indicate that our ED method that leverages MS loss is an effective way to use pseudo labels. For Pse-Augment, adding source domain entity labels to the target task leads to the negative transfer (NT) problem Zhang et al. (2020). For Pse-Classifier, the pipeline suffers from error propagation, where errors caused by the classifier can severely affect the performance of target entity predictions.
Figure 3: Davies-Bouldin index criterion of clusters. For baseline and ED-concerned settings, pseudo entities are included and viewed in the same cluster as Disease.
How is the Representation Enhanced? To investigate the effect of our proposed methods, we project entity representations (i.e., averaged hidden states of entity tokens in input texts) into a two-dimensional space using t-SNE van der Maaten and Hinton (2008). We show the visualization of target task BC5CDR in Appendix F. For better comparison, we adopt the Davies-Bouldin index (DB) Davies and Bouldin (1979) as the criterion. The lower the DB is, the better the data points are clustered.
With our ED method, the model effectively learns to disperse the representations of chemical entities and pseudo labeled entities. Therefore, it becomes easier for our model to assign a negative label to a source domain entity by measuring similarities between its representations and the target domain entity representations. Furthermore, our EG method also plays a vital role in the formation of feature space of the target domain. In the projection, chemical entities are more clearly clustered and disease entities are dispersed from chemical entities, which makes chemical entities easier to be extracted. For a more precise comparison, Figure 3 reports the DB index. Our methods achieve lower DB compared to the baselines, which indicates that the improved representations of the target domain are learned with our methods.
5 Related Work
Biomedical and Chemical Entity Extraction.
Entity extraction is a primary step in facilitating scientific discovery. Previous biomedical entity extraction methods can be categorized into several classes, including domain adaptive pretraining Labrak et al. (2023), boundary denoising diffusion model Shen et al. (2023a), question answering-based classification Arora and Park (2023), Cocke-Younger-Kasami (CYK) algorithm Corro (2023), in-context learning Chen et al. (2023), synthetic data Khandelwal et al. (2022); Chen et al. (2022); Hiebel et al. (2023), and prototype learning Cao et al. (2023).
Although there’s a shared corpus between the biomedical and chemical domains, entity extraction in the chemical domain remains underexplored. The chemical entity extraction task is usually viewed as an auxiliary task for biomedical named entity recognition Phan et al. (2021); Kocaman and Talby (2021); Luo et al. (2022); Lee et al. (2023). Similar to Nguyen et al. (2022), our paper differs from previous papers by viewing the chemical domain as an independent subject. Previous methods try to address this task by distant-supervision Wang et al. (2021b) or span-representation learning Nguyen et al. (2023). On the contrary, given the shared corpus between the biomedical and chemical domains, we leverage the large labeled data in the biomedical domain through transfer learning.
Transfer Learning for Name Tagging . Transfer learning is an effective method to address low-resource named entity recognition tasks Lee et al. (2018); Cao et al. (2018) and has shown its effectiveness in the biomedical domain Peng et al. (2019). Prior work has explored the role of continual pretraining on the target domain data Gururangan et al. (2020); Liu et al. (2021b). However, Mahapatra et al. (2022) argues that continual pretraining is inefficient regarding computational resources. In contrast to domain-adaptive pretraining, we aim to improve the representation of entities by projecting the source and target entities into separate regions of feature space. Inspired by the success of incorporating external knowledge for biomedical information extraction Zhang et al. (2021); Banerjee et al. (2021), we use biomedical events to augment the representations. To separate the potentially false positive examples in the target domain, we introduce pseudo-labels. Previous papers have adopted pseudo-labels in cross-lingual named entity recognition Zhou et al. (2023); Ma et al. (2023). However, they aim to align the entities in the source and target language instead of separating source entities from target entities. Compared to our method, Zhou et al. (2023) has a different contrastive objective, which aims to separate different entity types, rather than the entities from the two domains.
6 Conclusions
We proposed a named entity recognition task for transferring knowledge from the biomedical domain to the chemical domain. Our core idea is to train a shared feature space between the two domains to facilitate the knowledge transfer, and, at the same time, to project the source and target entities into separate regions of the feature space to reduce the false negative rate. We achieve this in a few steps. We begin by enriching the source feature space with information about events, then train a name tagging model to cluster similar entities into groups. We then use the trained model to label the entities that may belong to the source domain, and use these entities in a multi similarity loss function to achieve our goal. Our experiments across three source and three target datasets signify the effectiveness of our method.
7 Limitations
Our method partly relies on external knowledge. Therefore, the quality of external knowledge significantly influences the effect of our method. Especially when human annotations are unavailable, the performance of automatic annotators, typically neural networks, is an important factor to consider.
In this paper, we propose a framework that incorporates external knowledge during training. For instance, the compression function in §3.1 and templates in Section B.2 can be altered. Besides, such designs require prior knowledge of the source domain.
Our method leverages a contrastive learning strategy. However, the training algorithm doesn’t fully utilize GPU resources, leading to training inefficiencies.
Appendix A Details of ChatGPT evaluation on CHEMDNER
To evaluate ChatGPT’s few-shot ability of name tagging in the chemical domain, we elaborately select three name tagging examples in the CHEMDNER dataset that include all types of entities. An instructional description of the CHEMDNER task and examples constitute the prompt used to instruct the prediction. Configuration of prompts is shown in Table 5. The precision, recall, and macro-F1 scores on the CHEMDNER test set are $11.09\%$, $35.37\%$, and $16.88\%$ respectively. Laskar et al. (2023) reports the ChatGPT’s name tagging performance on general domain name tagging task-WNUT 17 dataset Derczynski et al. (2017) with precision: $18.03\%$, recall: $56.16\%$, and F1: $27.03\%$. This highlights ChatGPT’s weaker performance in solving chemical domain name tagging tasks.
Role
Prompt
System
You are an expert of chemical named entity recognition tasks
User
Description: In this task, you are given a small paragraph of a PubMed article, and your task is to identify all the named entities (particular chemical related entity) from the given input and also provide type of the each entity according to structure-associated chemical entity mention classes (ABBREVIATION, IDENTIFIER, FORMULA, SYSTEMATIC, MULTIPLE, TRIVIAL, FAMILY). Specifically, the paragraph are given with seperate tokens and you need to list all the chemical named entities in order and also tag their types. Generate the output in this format: entity1 <type_of_entity1>, entity2 <type_of_entity2>.
Examples:
Input: In situ C-C bond cleavage of vicinal diol following by the lactolisation resulted from separated treatment of Arjunolic acid ( 1 ) , 24-hydroxytormentic acid ( 2 ) and 3-O-$\beta$-D-glucopyranosylsitosterol ( 3 ) with sodium periodate and silica gel in dried THF according to the strategic position of hydroxyl functions in the molecule .
Output: C-C <FORMULA>, vicinal diol <FAMILY>, Arjunolic acid <TRIVIAL>, 24-hydroxytormentic acid <SYSTEMATIC>, 3-O-$\beta$-D-glucopyranosylsitosterol <SYSTEMATIC>, sodium periodate <SYSTEMATIC>, silica gel <TRIVIAL>, THF <ABBREVIATION>, hydroxyl <SYSTEMATIC>
Input: Structural studies using LC/MS/MS analysis and ( 1 ) H NMR spectroscopy showed the formation of a glycosidic bond between the primary hydroxyl group of RVX-208 and glucuronic acid .
Output: ( 1 ) H <FORMULA>, primary hydroxyl <FAMILY>, RVX-208 <IDENTIFIER>, glucuronic acid <TRIVIAL>
Input: The lystabactins are composed of serine ( Ser ) , asparagine ( Asn ) , two formylated/hydroxylated ornithines ( FOHOrn ) , dihydroxy benzoic acid ( Dhb ) , and a very unusual nonproteinogenic amino acid , 4,8-diamino-3-hydroxyoctanoic acid ( LySta ) .
Output: lystabactins <FAMILY>, serine <TRIVIAL>, Ser <FORMULA>, asparagine <TRIVIAL>, Asn <FORMULA>, formylated/hydroxylated ornithines <MULTIPLE>, FOHOrn <ABBREVIATION>, dihydroxy benzoic acid <SYSTEMATIC>, Dhb <ABBREVIATION>, 4,8-diamino-3-hydroxyoctanoic acid <SYSTEMATIC>, LySta <ABBREVIATION>
Please continue:
Input: %s
Output:
Table 5: Prompts of ChatGPT evaluation. %s is to be replaced by test context.
Appendix B Details of Event Embedding Construction.
B.1 Concatenation based Event Embedding
An overview of concatenation based event embedding strategy is shown in Figure 4. Each type of annotated event contains a trigger and various arguments, including theme, cause, product, and site. We prepare raw texts of each argument and encode them into argument embeddings. To illustrate the formation of raw texts, let’s consider an example. Imagine a gene named “IL-1ra” which is associated with two event annotations, “M1, Negation, E9” and “E9, Binding:forms a complex, Theme:IL-1ra, Theme2:Type I IL-1R”. Raw text for “event_type” comprises the event name, “M” label, and trigger. The example’s raw text should be “Binding (Negation): forms a complex”. Raw text typically is the corresponding entity itself for the rest of the arguments. However, for the focusing entity, “IL-1ra” for instance, raw text is specified as “self”, deriving “IL-1ra (self)” in this case. There are several corner cases to tackle with:
Nested Event.
As mentioned in §3.1, for the nested event, we first recursively compose the event embedding for it and compress it to the same length as partial embedding. To achieve this, let’s consider the nested event embedding as $e$. We then implement the compression function by averaging several successive elements:
$f(e)=\left[\frac{1}{5}\sum_{k=0}^{4}e_{k},\frac{1}{5}\sum_{k=5}^{9}e_{k},...,\frac{1}{5}\sum_{k=5i}^{5i+4}e_{k},...\right]^{T},$ (12)
Full embedding has $768\times 5$ dimensions, and we average every $5$ element and concatenate the values into a $768$-dimension embedding.
Padding
It is necessary that some arguments do not apply to some events or miss in annotation. A padding scheme is necessary for missing arguments, and we choose to fill in random partial embedding with the same length sampled from Gaussian distribution with the same mean and covariance as all other encoded partial embeddings.
B.2 Sentence-Encoder based Event Embedding
The prompt we use to instruct ChatGPT to generate explanatory templates for events is: give a one-sentence definition of biomedical event type XXX with arguments XXX, XXX…. For instance, we generate a template for the Phosphorylation event with give a one-sentence definition of biomedical event type Phosphorylation with arguments Trigger, Theme:Molecule, Cause:Molecule, Site:Simple chemical. The full list of the used templates is shown in Table 9, 10, 11, 12, and 13.
Appendix C Experiment Details
We select bert-base-uncased version of BERT model as our backbone model, and we only train 0.817% of parameters (894,528) of the entire model using transformer-adapter utils Pfeiffer et al. (2020). For source task pretraining, we use a batch size of 64; while for target task fine-tuning, we use a batch size of 16, considering the relatively small training set. We use AdamW optimizer and an initial learning rate of $41e-$ 4 for pretraining and finetuning. To fully train our model, we first train 80 epochs and then stop when f1 scores on the held-out validation set fail to update the best score for at least 20 epochs in a row. For hyperparameters, we tune balance factor $\lambda_{\mathbb{S}}$ from scale $\{0.10,0.15,0.20,0.25,0.30\}$ and tune $\lambda_{\mathbb{T}}$ from scale $\{0.6,0.8,1.0,1.2,1.4\}$. Due to the limitation of computational resources, we select $\epsilon$ and $\gamma$ as $0.1$ and $0.5$ respectively, and $\alpha$, $\beta$, $\rho$, $\tau$ as $4.0$, $3.0$, $8.0$, $6.0$ respectively considering the ratio of number of positive and negative pairs in contrastive learning.
Our models are trained on 4 Nvidia RTX 2080Ti GPUs in a data parallel fashion. Source task pretraining with contrastive learning takes around 5 hours, while target task finetuning with contrastive learning takes around 30 minutes.
Appendix D Full Experiment Results
More experiment results are reported in Table 8.
Target Tasks
CHEMD BC5CDR DrugProt
Direct Transfer
46.18
78.55
68.17
EG(MS)-Only
46.08
78.87
68.00
EG(concat)-Only
47.03
78.89
67.49
EG(sentEnc)-Only
46.91
79.76
67.93
Table 6: F1(%) scores on three target tasks. Performance of our EG method using vanilla MS loss without external knowledge is reported as EG(MS)-Only.
Appendix E Ablation Experiment Details
E.1 Generalization Ability
Event Annotator
Target Tasks
CHEMD BC5CDR DrugProt
Gold-std
Concat
47.03
78.89
67.49
SentEnc
46.91 79.76
67.93
Auto-sys
Concat
46.06
79.49 68.31
SentEnc
46.11
79.39
68.04
Table 7: F1(%) scores of our proposed EG methods based on human/machine annotated events. We highlight better scores between Gold-std and Auto-sys annotators under each setting with underlines.
To alleviate the reliance on gold-standard event annotations, which may be hard to obtain, we generate the event annotations using DeepEventMine Trieu et al. (2020). Table 7 reports the performance of the EG methods. We see that the performance is comparable to that of the gold-standard annotations. We also observe that in the DrugProt dataset, the performance with automatic annotations is better than that of the gold-standard annotations, suggesting that the human annotations are low quality and noisy.
E.2 External Knowledge
Table 6 reports the performance of our EG method without external knowledge (i.e., event annotations), where simple MS loss replaces our RMS loss. The performance over three target tasks mirrors the Direct Transfer setting, suggesting that the vanilla MS objective has minimal impact and the main improvement stems from the auxiliary data extracted from the event mentions.
E.3 Pseudo Label Usage
Pse-Augment
We augment annotations of target tasks with pseudo entities within the target corpus labeled as “Out of Distribution (OOD)” entities. Then, the model is trained with a single cross-entropy loss.
Pse-Classifer
We first use a pretrained model as a classifier separating pseudo entities and gold-standard annotated entities. We then predict with the former directly finetuned model and filter out entities labeled “OOD” by the classifier.
Appendix F Visualization
The t-SNE projection visualization of BC5CDR test entity representations is shown in Figure 5.
Figure 5: t-SNE visualization of entities in the test corpus of BC5CDR. Pseudo is labeled by model pretrained on source task, Disease and Chemical are gold-standard annotations. BERT represents vanilla BERT model without pretraining or finetuning, and all the settings are same as Main Results.
Appendix G Scientific Artifacts
We list the license used in this paper: PathwayCuration (CC BY-SA 3.0), InfectiousDisease (CC BY-SA 3.0), CancerGenetics (CC BY-SA 3.0), CHEMDNER (CC BY 4.0), BC5CDR (CC BY 4.0), DrugProt (CC BY 4.0), Huggingface Transformers (Apache License 2.0), SapBERT (apache-2.0), S-PubMedBert-MS-MARCO-SCIFACT (apache-2.0), OpenAI (Terms of use666openai.com/policies/terms-of-use). We follow the intended use of all the mentioned existing artifacts in this paper.
Evaluate Datasets
CHEMDNER BC5CDR DrugProt
Metrics
Precision
Recall
F1
Precision
Recall
F1
Precision
Recall
F1
Few-shot (BERT)
42.73
51.69
46.77
72.44
85.86
78.51
63.80
67.42
65.49
Pathway Curation
Direct Transfer
42.51
49.70
45.82
66.79
88.79
76.22
60.94
68.70
64.52
EG(concat)-Only
44.75
51.78
48.00
71.49
87.04
78.45
64.45
66.39
65.31
EG(sentEnc)-Only
42.80
50.43
46.25
70.91
86.97
78.09
64.36
67.81
66.03
ED-Only
46.97 52.30 49.48
74.71
83.96
79.06
67.05
67.13
67.08
EG(concat)+ED
43.44
51.54
47.11
72.34
86.09
78.61
66.59
66.13
66.32
EG(sentEnc)+ED
43.34
51.73
47.16
75.66
86.14
80.55
66.15
68.32
67.21
Infectious Diseases
Direct Transfer
41.57
48.30
44.65
74.92
82.51
78.53
61.96
67.81
64.75
EG(concat)-Only
45.26
50.65
47.77
72.28
87.41
79.12
63.76
65.25
64.43
EG(sentEnc)-Only
47.33
50.41
48.80
74.58
86.37
80.04
64.83
63.60
64.20
ED-Only
41.07
46.93
43.78
74.60
85.85
79.80
64.25
69.75
66.63
EG(concat)+ED
43.37
50.26
46.43
73.83
85.48
79.23
62.31
66.87
64.47
EG(sentEnc)+ED
41.50
52.19
46.18
74.70
85.95
79.93
65.06
69.11
67.02
Cancer Genetics
Direct Transfer
45.22
51.80
48.27
72.37
86.56
78.82
62.94
69.61
66.07
EG(concat)-Only
40.59
53.19
46.01
71.44
87.22
78.53
65.31
66.67
65.98
EG(sentEnc)-Only
41.54
52.36
46.33
72.68
86.09
78.82
66.99
65.71
66.30
ED-Only
47.06
50.58
48.75 75.08
86.96
80.57
66.26
73.52 69.68
EG(concat)+ED
45.16
51.72
48.16
73.37
85.06
78.76
65.60
67.67
66.64
EG(sentEnc)+ED
41.89
50.17
45.61
73.81
86.03
79.45
65.59
68.47
66.99
Table 8: Full evaluation results. Pathway Curation, Infectious Diseases and Cancer Genetics are set to be the source domain respectively. Experiment settings are the same as main results reported in Table 3.
Type
Template
Conversion
A specific trigger <Trigger> causes the transformation of a molecule <Theme> into another molecule <Product>.
Phosphorylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the addition of a phosphate group at a particular site <Site>, facilitated by another molecule <Cause>.
Dephosphorylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a phosphate group removed from a particular site <Site>, facilitated by another molecule <Cause>.
Acetylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of an acetyl group at a particular site <Site>, catalyzed by another molecule <Cause>.
Deacetylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has an acetyl group removed from a particular site <Site>, facilitated by another molecule <Cause>.
Ubiquitination
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the attachment of one or more ubiquitin molecules at a particular site, facilitated by another molecule <Cause>, often involving a simple chemical group <Site> as the site of attachment.
Deubiquitination
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has ubiquitin molecules removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
Hydroxylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of a hydroxyl group at a particular site, catalyzed by another molecule <Cause> involving a simple chemical group <Site> as the site of attachment.
Dehydroxylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a hydroxyl group removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
Methylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of a methyl group at a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of attachment.
Demethylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a methyl group removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
Localization
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is directed to or away from a particular cellular component <AtLoc><FromLoc><ToLoc> or subcellular location within the cell.
Transport
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is moved or conveyed to or away from a particular cellular component <FromLoc><ToLoc> or subcellular location within the cell.
Gene Expression
Indicated by the given trigger <Trigger>, genetic information from a gene <Theme> is used to produce a functional gene product, such as RNA or protein.
Transcription
Indicated by the given trigger <Trigger>, genetic information from a gene <Theme> is transcribed into RNA, usually messenger RNA (mRNA), by RNA polymerase.
Translation
Indicated by the given trigger <Trigger>, the genetic information carried by mRNA <Theme> is used to synthesize a protein by ribosomes in the cell.
Degradation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes breakdown or degradation into smaller components.
Binding
Indicated by the given trigger <Trigger>, a specific molecule <Theme> interacts and forms a complex with another molecule(s) resulting in the product of a molecular complex <Product>.
Table 9: Templates for PathwayCuration Dataset.
Type
Template
Binding
Indicated by the given trigger <Trigger>, a specific molecule <Theme> interacts and forms a complex with another molecule(s) resulting in the product of a molecular complex <Product>.
Dissociation
Indicated by the given trigger <Trigger>, a specific complex <Theme> breaks apart, resulting in the release of individual molecules <Product> as products.
Regulation
Indicated by the given trigger <Trigger>, an entity <Theme> is controlled or influenced by another entity <Cause> to achieve a specific biological effect or outcome.
Positive Regulation
Indicated by the given trigger <Trigger>, an entity <Theme> is promoted or enhanced by another entity <Cause> to achieve a specific biological effect or outcome.
Activation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is stimulated or facilitated by another entity <Cause> to increase its activity, function, or biological effect.
Negative Regulation
Indicated by the given trigger <Trigger>, an entity <Theme> is inhibited or suppressed by another entity <Cause> to achieve a specific biological effect or outcome.
Inactivation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is deactivated or rendered inactive by another entity <Cause>, leading to a reduction or cessation of its biological function.
Pathway
Indicated by the given trigger <Trigger>, involving one or more molecules <Participant> that collaborate to accomplish a specific biological function or response.
Table 10: Continuation of templates for PathwayCuration Dataset.
Type
Template
Gene Expression
Indicated by the given trigger <Trigger>, a specific protein or a group of genes <Theme> are involved in the transcription and translation of genetic information to produce functional gene products, such as RNA or protein.
Transcription
Indicated by the given trigger <Trigger>, a specific protein or a group of genes <Theme> are involved in the synthesis of RNA from DNA template by RNA polymerase.
Protein Catabolism
Indicated by the given trigger <Trigger>, a specific protein <Theme> is broken down or degraded into smaller peptide fragments or amino acids.
Phosphorylation
Indicated by the given trigger <Trigger>, a specific protein <Theme> undergoes the addition of a phosphate group at a particular site <Site>, resulting in the modification of the protein’s structure and function.
Localization
Indicated by the given trigger <Trigger>, a specific core entity <Theme> is directed to or away one location <AtLoc><ToLoc> within the cell or organism.
Binding
Indicated by the given trigger <Trigger>, a specific core entity <Theme> interacts and forms a connection with another entity <Site>, leading to the formation of a complex or association.
Regulation
Indicated by the given trigger <Trigger>, a specific core entity or event <Theme> is controlled or influenced by another core entity or event <Cause> through interactions at specific sites on molecules or entities <Site><CSite>, potentially resulting in modulation of biological processes.
Positive Regulation
Indicated by the given trigger <Trigger>, a specific core entity or event <Theme> is promoted or enhanced by another core entity or event <Cause> through interactions at specific sites on molecules or entities <Site><CSite>, potentially resulting in an increase in the intensity or rate of a biological process.
Negative Regulation
Indicated by the given trigger <Trigger>, a specific core entity or event <Theme> is inhibited or suppressed by another core entity or event <Cause> through interactions at specific sites on molecules or entities <Site><CSite>, potentially resulting in a decrease in the intensity or rate of a biological process.
Process
Indicated by the given trigger <Trigger>, involving a core entity that collaborates to accomplish a specific biological function or response.
Table 11: Templates for Infectious Diseases Dataset.
Type
Template
Development
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> undergoes progressive changes or growth, leading to the formation of a more complex and specialized structure or condition over time.
Blood Vessel Development
Indicated by the given trigger <Trigger>, blood vessels <Theme> undergo progressive changes or growth at a specific location <AtLoc>, leading to the formation and maturation of the vascular network.
Growth
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> undergoes an increase in size, quantity, or complexity over time.
Death
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> ceases to exhibit signs of life and undergoes irreversible loss of vital functions.
Cell Death
Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes a series of events leading to its own demise, often through programmed cell death or other mechanisms.
Breakdown
Indicated by the given trigger <Trigger>, a specific anatomical or pathological structure <Theme> disintegrates, decomposes, or undergoes degradation over time.
Cell Proliferation
Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes rapid and controlled replication or division, leading to an increase in the number of daughter cells.
Cell Division
Indicated by the given trigger <Trigger>, a specific cell <Theme> divides into two or more daughter cells through mitosis or meiosis.
Cell Differentiation
Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes changes in gene expression and morphology to become specialized and acquire distinct functions at a specific anatomical or pathological location <AtLoc>.
Remodeling
Indicated by the given trigger <Trigger>, a specific tissue <Theme> undergoes structural changes, reorganization, and modification in response to various stimuli or during growth and development.
Reproduction
Indicated by the given trigger <Trigger>, a specific organism <Theme> produces offspring through sexual or asexual means, leading to the continuation of the species.
Mutation
indicated by the given trigger <Trigger>, a specific gene, genome, or protein <Theme> undergoes a heritable change in its genetic sequence or structure at a particular site <Site>, potentially leading to alterations in its function or expression within a specific anatomical or pathological context <AtLoc>.
Carcinogenesis
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> undergoes a series of genetic and cellular changes at a specific anatomical or pathological location <AtLoc>, leading to the development of cancer.
Cell Transformation
Indicated by the given trigger <Trigger>, a specific cell <Theme> undergoes changes in its phenotype, function, or behavior at a specific anatomical or pathological location <AtLoc>, often associated with the acquisition of abnormal or cancerous characteristics.
Metastasis
Indicated by the given trigger <Trigger>, a specific anatomical or pathological entity <Theme> spreads and establishes secondary growths or lesions at a different anatomical or pathological location <ToLoc> from the primary tumor site.
Infection
Indicated by the given trigger <Trigger>, an organism <Participant> invades and establishes itself in a specific anatomical or pathological site <Theme>, leading to a disease condition.
Metabolism
Indicated by the given trigger <Trigger>, collective chemical reactions occur within an organism <Theme> involving the processing, transformation, and utilization of specific molecules to maintain cellular functions and energy requirements.
Synthesis
Indicated by the given trigger <Trigger>, a specific simple chemical <Theme> is produced or created through chemical reactions or biological processes.
Catabolism
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes chemical reactions or metabolic pathways to break down into simpler compounds, releasing energy in the process.
Amino Acid Catabolism
Indicated by the given trigger <Trigger>, a specific amino acid <Theme> is broken down through metabolic pathways, leading to the release of energy and the generation of byproducts like ammonia and carbon compounds.
Glycolysis
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes a series of enzymatic reactions, ultimately converting glucose into pyruvate and producing ATP and NADH as energy carriers.
Table 12: Templates for Cancer Genetics Dataset.
Type
Template
Glycosylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of carbohydrate molecules (glycans) to specific sites, typically on proteins or lipids, leading to the formation of glycoproteins or glycolipids with diverse biological functions.
Acetylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of an acetyl group at a particular site <Site>, catalyzed by another molecule <Cause>.
Deacetylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has an acetyl group removed from a particular site <Site>, facilitated by another molecule <Cause>.
Ubiquitination
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is modified by the attachment of one or more ubiquitin molecules at a particular site, facilitated by another molecule <Cause>, often involving a simple chemical group <Site> as the site of attachment.
Deubiquitination
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has ubiquitin molecules removed from a particular site, facilitated by another molecule <Cause> involving a simple chemical group <Site> as the site of removal.
Gene Expression
Indicated by the given trigger <Trigger>, a specific gene, genome, or protein <Theme> is activated, leading to the production of RNA or protein and subsequent biological functions.
Transcription
Indicated by the given trigger <Trigger>, genetic information from a specific gene, genome, or RNA molecule <Theme> is used as a template to produce complementary RNA (usually messenger RNA) by RNA polymerase.
Translation
Indicated by the given trigger <Trigger>, the genetic information carried by messenger RNA <Theme> is used to synthesize a protein by ribosomes in the cell.
Protein Processing
Indicated by the given trigger <Trigger>, the series of post-translational modifications, folding, and transportation of a specific gene, genome, or protein <Theme> in the cell to achieve its mature and functional form.
Phosphorylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> undergoes the addition of a phosphate group at a particular site <Site> within a protein domain or region, often regulating the molecule’s activity or function.
Dephosphorylation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> has a phosphate group removed from a particular site <Site> within a protein domain or region, often resulting in the modulation or termination of its activity or function.
DNA Methylation
Indicated by the given trigger <Trigger>, a specific gene or genome <Theme> has a methyl group added to a particular site <Site> within a protein domain or region, often resulting in the regulation of gene expression and epigenetic modifications.
DNA Demethylation
Indicated by the given trigger <Trigger>, a specific gene or genome <Theme> has a methyl group removed from a particular site <Site> within a protein domain or region, often resulting in the modulation of gene expression and epigenetic modifications.
Pathway
Indicated by the given trigger <Trigger>, involving a specific molecule <Participant> that collaborates to accomplish a specific biological function or response.
Binding
Indicated by the given trigger <Trigger>, a specific molecule <Theme> forms a physical association or interaction with another molecule or region <Site> on a protein or DNA, potentially leading to functional changes or regulatory effects.
Dissociation
Indicated by the given trigger <Trigger>, a specific molecule <Theme> separates or detaches from another molecule or region <Site> on a protein or DNA, leading to the termination or disruption of their interaction or complex formation.
Localization
Indicated by the given trigger <Trigger>, a specific molecule <Theme> is directed to or away from a particular cellular component <AtLoc><FromLoc><ToLoc> or subcellular location within the cell.
Regulation
Indicated by the given trigger <Trigger>, an entity <Theme> is controlled or influenced by another entity <Cause> to achieve a specific biological effect or outcome.
Positive Regulation
Indicated by the given trigger <Trigger>, an entity <Theme> is promoted or enhanced by another entity <Cause> to achieve a specific biological effect or outcome.
Negative Regulation
Indicated by the given trigger <Trigger>, an entity <Theme> is inhibited or suppressed by another entity <Cause> to achieve a specific biological effect or outcome.
Planned Process
Indicated by the given trigger <Trigger>, a specific entity <Theme> is involved or manipulated using an instrument <Instrument> for a predetermined outcome or purpose.
Table 13: Continuation of templates for Cancer Genetics Dataset.